OpenStudy (sodapop):

A group of people were given a personality test to determine if they were type A or Type B. The results are shown below. Male Female Type A 65 85 Type B 38 12

OpenStudy (sodapop):

Im writing the answers

OpenStudy (sodapop):

Compare P(Male or type B) with P(male / Type B) P(Male or Type B) > P(Male /Type B) P(male or type B) = P(Male / type B) P(Male or Type B) < P(Male/ type B) There is not enough information

OpenStudy (jim766):

to find the P(male or Type B) P(male) + P(type B) - P(male and Type B) do you know how to find the probability of someone in this group being male ?

OpenStudy (sodapop):

No please expalain

OpenStudy (jim766):

p(male) = Total males/total people add up the totals in your table

OpenStudy (jim766):

how many males in the survey?

OpenStudy (jim766):

65 + 38 =

OpenStudy (sodapop):

103

OpenStudy (jim766):

how many people in the survey?

OpenStudy (jim766):

Males and females

OpenStudy (sodapop):

200

OpenStudy (jim766):

right so p(male) = 103/200 Now do the same thing to find P(type B)

OpenStudy (jim766):

how many people were type B?

OpenStudy (jim766):

males and females

OpenStudy (sodapop):

50

OpenStudy (jim766):

right, and what was total # of people?

OpenStudy (sodapop):

200

OpenStudy (sodapop):

So 200/50 ?

OpenStudy (jim766):

so P(type B) = 50/200

OpenStudy (sodapop):

Which equals 0.25

OpenStudy (jim766):

p(male or type B) = 103/200 + 50/200 - p(male and type B) how many people in the survey were male and type b?

OpenStudy (jim766):

look at the table..

OpenStudy (sodapop):

38 were male

OpenStudy (jim766):

p(male or type B) = 103/200 + 50/200 - 38/200

OpenStudy (jim766):

the denominators are all the same, so add up the numerators and divide by 200

OpenStudy (sodapop):

.955

OpenStudy (jim766):

103+50-38 = 115/200

OpenStudy (jim766):

or ,575

OpenStudy (sodapop):

Alright I forgot to subtract

OpenStudy (jim766):

P ( male / Type B) means probability that someone is male given they are type B the slash means given

OpenStudy (jim766):

in this case we dont care about everyone, we just care about the people who were type B

OpenStudy (sodapop):

Alright

OpenStudy (jim766):

look at the type B row males/ total of the type B people

OpenStudy (jim766):

P ( male / Type B) = 38/ 50

OpenStudy (sodapop):

38/50

OpenStudy (sodapop):

Ok

OpenStudy (jim766):

or .76

OpenStudy (sodapop):

So would the answer be C ?

OpenStudy (jim766):

which is bigger?

OpenStudy (jim766):

Right!

OpenStudy (anonymous):

agggg Thanks sooo much <333

OpenStudy (pink.1d):

@malice ;)